OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

OpenStudy (anonymous):

Two Point form is given by: \[\huge \color{blue}{\frac{y - y_1}{y_2 - y_1} = \frac{x- x_1}{x_2-x_1}}\]

OpenStudy (anonymous):

Appreciate it Bro